OpenStudy (anonymous):

What is the number of real solutions? –11x^2 = x + 11

OpenStudy (anonymous):

no one is helping?? that s a shame!

OpenStudy (anonymous):

Can you?

OpenStudy (anonymous):

Find discriminant and then use quadratic formula to find the roots of equaton

OpenStudy (anonymous):

Do you still need help with this one?

OpenStudy (anonymous):

start with \[11x^2+x+11=0\] then use \(b^2-4ac\) and you will see that it is negative, so no real zeros

OpenStudy (anonymous):

So that means it has no real solutions?

OpenStudy (anonymous):

since \[b^2-4ac=1-4\times 11\times 11<0\] (aka negative) there will be no real solutions the quadratic formula has a \[\sqrt{b^2-4ac}\] in it, so if the inside part is negative, no real solutions

OpenStudy (anonymous):

okay , thank you. Can you also tell me how to identify 3i on a graph?

OpenStudy (anonymous):

if it is the graph of the complex plain, think of it as the point \((0,3)\)

OpenStudy (anonymous):



OpenStudy (anonymous):

okay, that makes since. Thank you

OpenStudy (anonymous):

yw